### v1

# How should AI decisions be explained?

Abstract.
This paper investigates the relationship between law and eXplainable Artificial Intelligence (XAI). While there is much discussion about the AI Act, for which the trilogue of the European Parliament, Council and Commission recently concluded, other areas of law seem underexplored. This paper focuses on European (and in part German) law, although with international concepts and regulations such as fiduciary plausibility checks, the General Data Protection Regulation (GDPR), and product safety and liability. Based on XAI-taxonomies, requirements for XAI-methods are derived from each of the legal bases, resulting in the conclusion that each legal basis requires different XAI properties and that the current state of the art does not fulfill these to full satisfaction, especially regarding the correctness (sometimes called fidelity) and confidence estimates of XAI-methods.
Law, EU, European Law, German Law, eXplainable Artificial Intelligence (XAI), XAI, Interpretable ML, norms, standards, regulation, safe ML
copyright: noneccs: Social and professional topics Governmental regulationsccs: Social and professional topics Management of computing and information systemsccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligence
1. Introduction
The explainability of AI systems is crucial for their development and deployment in sensitive domains. Therefore, the AI Act draft of the European Union (EU) mandates that high-risk AI systems ensure “sufficient transparency” for users to interpret and utilize the system’s results appropriately. However, a challenge lies in the vagueness of these transparency requirements and what exactly constitutes “sufficient transparency”. While more details on this “sufficient transparency” will hopefully be specified in standards such as the ISO/IEC CD TS 6254, other legal bases for the use of eXplainable Artificial Intelligence (XAI) already exist. This paper reviews the legal requirements for explainability of AI systems based on European (and in part German) law, focusing on three legal domains and four application areas. Therefore, our contributions can be summarized as follows:
  • We present an extended list of XAI properties which can be used to formulate requirements.
  • We describe legal bases for the use of XAI and derive requirements such methods need to fulfill in each of the considered areas of law.
  • We point out open research problems and a possible way forward for the field of XAI by embracing interdisciplinarity.
For this, we first provide a brief overview of common XAI terminology in the next section and establish the choice of areas of law. After the description of related works in Chapter 2, the XAI properties used in this paper are described in Chapter 3, with the distinction of method-specific properties and process-specific properties. In Chapter 4, each of the considered areas of law is introduced, accompanied by the formulation of the task for XAI in this area, from which the XAI requirements are derived. The paper is closed with a discussion in Chapter 5 and a summary in Chapter 6. The paper structure and requirements for XAI are also replicated in short in Table 1 in Chapter 6.
$1.$ Taxonomy
We start by summarizing the taxonomy of explainability methods for AI briefly. For a full introduction we point the readers to (Burkart and Huber, 2021; Molnar, 2022). On the highest level, explanations can be divided into local and global ones. Local explanations explain a particular decision (e.g., the loan was rejected due to a low income) while global explanations reveal how decisions are made in general (e.g., every applicant with income above 100k is given a loan). They are introduced in more detail below. It needs to be noted that most XAI-methods provide non-causal explanations, although special causal methods exist.
1.1.1. Local Explanations
Feature attributions are widely used to provide local explanations by quantifying the contribution of specific features to a decision. For example, in the loan granting scenario, feature attribution could reveal that a short length of employment had the highest negative impact on the rejection, while an income of 75k was the strongest positive feature. Methods like SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017) and Linear Interpretable Model-Agnostic Explanations (LIME) (Ribeiro et al., 2016) are commonly used for computing feature attributions. Saliency maps, on the other hand, highlight important regions in images that contribute to specific decisions. They can be created using methods like SHAP or LIME, although usually methods specifically for neural networks—such as Grad-CAM (Selvaraju et al., 2020)—are used. Counterfactual explanations are a type of local explanation that provide insight into decision-making without revealing the detailed functioning of the AI model (Wachter et al., 2018). These explanations show a change in input features necessary to alter the outcome. For instance, a counterfactual explanation for the loan example could be “If you had worked at least five years, you would have been granted the loan.” Data attribution or sample-based explanations involve identifying influential samples from the training set that contributed to a decision. In the loan example, this could be interpreted as “Your loan was rejected because similar applicants from the training data defaulted on a loan,” potentially accompanied by such similar applicants.
1.1.2. Global Explanations
Global explanations provide information about the workings of a model across its entire data range. The simplest form of global explainability is the usage of so-called whitebox models such as decision trees. These models provide out-of-the-box insights into the decision making process and, hence, do not need any specific post-hoc explainability—given that they are small enough in size to be understandable by humans. Providing global explanations for black-box models is more difficult. Apart from partial dependence plots (PDPs), surrogate models provide a way of opening the black box to some degree. For this, either a black-box model is transformed into a white-box one (Nguyen et al., 2020) or a white-box (surrogate) model is fitted on the output of the black-box model that shall be explained, thus showing how similar decisions could be made. As a further way to explain neural networks, especially in computer-vision tasks, concept-based explanations can be used. They usually provide an account of which concepts, e.g., certain patterns like stripes or dots, are especially relevant for the classification of single images (local concept-based explanations) or for the class or model under consideration in general (global concept-based explanations) (Kim et al., 2018). Other methods produce global feature attribution explanations (Covert et al., 2020), highlighting features with high predictive power.
1.2. Considered Laws
This paper focuses on EU and German laws, although with a wide span of possible stakeholders and reasons for the use of XAI (see Chapter 4). While there is uncertainty about what XAI can achieve, the considered areas of law—company law, the General Data Protection Regulation (GDPR), product safety and liability for defect products—are among the most discussed and show the different legal questions XAI could possibly answer. Additionally, these areas bear similarity across many different jurisdictions, as the ISION-rule (as used in Chapter 4.1) serves as a specification of fiduciary duties, which are applicable to the directors and officers of nearly every corporation worldwide. Moreover, the ISION-rule is closely connected to § 141(e) of the Delaware General Corporation Law (Title 8, Chapter 1 of the Delaware Code) in the US, addressing the reliance on expert opinions by board directors. The GDPR (Chapter 4.2) was introduced as a de-facto international standard, coining the term “Brussels-Effect” for EU laws that are adhered to even outside the EU. Product safety and liability (Chapter 4.3) are harmonized within the EU. This paper does not intend to provide an exhaustive list of legal bases for the use of XAI, but aims to provide a starting point towards such a list and shows how laws can inform the development of XAI techniques.
2. Related Work
To date, not many works have concerned themselves with the link between law and XAI, apart from works around the AI Act and the GDPR. In the following, the related works and the understanding of transparency in law are summarized. A broader discussion around the need for algorithmic transparency was started by the provisions of the GDPR, especially about whether a “right to explanation” exists (Wachter et al., 2017) and how much information would be necessary to fulfill such a right. Wachter et al. (2018) devised counterfactual explanations to give users actionable advise while keeping the exact workings of an Machine Learning (ML) model a secret (see Section 1.1.1). Vorras and Mitrou (2021) argue for a “right to explanation” with the implication of a “qualified transparency mechanism”, i.e., adapting explanations to the target audience and “considering the technical limitations of AI interpretability, the underlying legal and regulatory barriers, as well as the legitimate expectations of stakeholders”. A similar sentiment can be found in (Błażej Kuźniacki, Marco Almada, Kamil Tyliński, Łukasz Górski, Beata Winogradska, Reza Zeldenrust, 2022), where Kuzniacki et al. argue for the construction of a global regulatory framework for the use of XAI in tax law. They additionally recommend that “tax lawyers and computer scientists should work together to help reach a consensus on legal and technical solutions that secure XAI’s place in tax law”. While our work is not based on tax law, it follows a similar idea as it aims to combine the information from legal experts and computer scientists to make progress towards a structured use of XAI for legal purposes. As of now, a gap between the legal and technical understanding can be discerned in many works, as echoed by Gyevnar et al. (2023), who argue that this difference in understanding can also be seen in the EU AI Act. They note that there seem to be multiple understandings of transparency and that the transparency mandated in the EU AI Act is different from purely technical transparency, which is the aim of most works in XAI. Also concerned with the different interpretations of transparency in the AI Act are Madalina Busuioc, Deirdre Curtin, Marco Almada (2022). They stress that the explainability of AI systems, i.e., the presentation of specific—mediated—details of their workings, is different from the disclosure of the model itself, where only the second option allows for the unbiased scrutiny of such systems. Such explanations and thus, mediated information, could be used maliciously, as shown by Zhou and Joachims (2023). Summarizing, despite some works in that direction, the use of XAI for legal purposes is still unclear.
3. Properties of XAI-Methods (possibly) relevant for their legal use
To be able to formulate requirements for XAI systems, a taxonomy of possibly necessary properties of such systems needs to be defined. In this work, we primarily rely on the 12 Co-properties introduced by Nauta et al. (2023), which were derived from scientific literature published until the end of 2020. This comprehensive set of properties is described below. Additional to these XAI-method specific properties, further ones are formulated in Section 3.2 for the process in which XAI is used. A short overview of the used properties can be found in Table 2 in the appendix.
3.1. 12 Co-Properties of Nauta
Correctness describes whether—or to what extent—an explanation conforms to the ML model the explanation was created for. While this seems like a very basic feature of any explanation method, this property is not necessarily fulfilled, for example for sample-based methods like LIME and SHAP. Completeness addresses how much of the explanandum—either a model itself or a prediction—is captured within an explanation. The extreme cases would be to either present the entire ML model with its mathematical operations or to just highlight one input feature as the main cause of a specific outcome. White-box models are by definition complete explanations of themselves. Consistency demands that identical inputs provide identical explanations, as some explanation methods are not deterministic in their behavior—especially sample-based methods. Inconsistent explanations might also result from the hyperparameter choices of specific methods (Tritscher et al., 2023). Continuity poses some similarity to Consistency, as it demands that small changes in the input or output of the model should also only result in small changes in the explanation. This can be illustrated for saliency maps for images: If an image is overlaid with noise, but the prediction of the model does not change, the explanation should also only change slightly. Contrastivity describes the discriminativeness of an explanation, as explanations are supposed to show why a decision was made in contrast to a different outcome that was not chosen, as argued by Miller (2019). Covariate Complexity shows how complex features and their interactions in an explanation are. The simplest way to go about this would be to use the input features and ignore all feature interactions (and thus possibly lose out on Correctness), whereas for example simplified feature interactions such as linear or monotonous interactions could be one step more complex. Compactness denotes the overall size of an explanation, as too large explanations often overwhelm users, whereas too small explanations provide less information about a decision or the model behavior (Hoffman et al., 2023). Compositionality describes how an explanation is presented, especially its format, organization, and structure. Different ways of presenting explanations can improve the way the explanation is received, for example by providing higher-level information (Alvarez-Melis and Jaakkola, 2018) or different terminology (Swartout and Moore, 1993). Confidence estimates can be used in two different ways to assist in the use of XAI—either to show the confidence of the model in the original prediction as part of the explanation or in denoting the confidence that a given explanation is correct and thus, conforms to the decision/model at hand. Context of an explanation, especially of who the explainee is, could help to adapt an explanation to specific user needs, thus improving its practical use. One of the earlier mentions of context can be found in (Miller et al., 2017) which even described the development of XAI methods only conforming to the needs of AI engineers and scientists as “inmates running the asylum.” Coherence addresses whether an explanation conforms to prior knowledge—to that of an explainee, general consensus or other beliefs—thus making it easier to understand. While humans often use their prior knowledge to evaluate whether they deem an explanations as trustworthy or correct, Coherence is distinctly different from Correctness as it is described above, as an explanation can be correct (i.e., conform to the logic of a model) without conforming to prior beliefs. Controllability describes how and to what extent an explanation can be interacted with. Also called interactivity, it seems as this can improve the understanding of an explanation (Cheng et al., 2019) as well and also helps to address Context, as users could adapt the explanation to their needs as they see fit. It should be noted that there is currently no consensus on the best way to evaluate the properties discussed, as different evaluation methods (with many listed in (Nauta et al., 2023)) may contradict each other (Tomsett et al., 2019). Therefore, this paper will not attempt to formulate specific requirements for validating XAI systems, such as through specific technical means or user studies.
3.2. Process Properties
The aforementioned properties are mainly formulated for XAI-methods themselves, while for legal purposes, requirements could also be formulated for properties of the implementation or of the process of creating explanations. These are discussed in the following.
Since most current XAI methods do not guarantee Correctness, using multiple methods can help, e.g., to spot errors in an application. Given their Consilience, this can also be an indication of Correctness. However, if explanations do not align, there is no established way to resolve such situations. The Computations in real-time systems should also be considered. In time-sensitive scenarios, explanations need to be provided from the start, while in other cases, longer runtimes (e.g., hours of runtime on a GPU-server) may be feasible. The Coverage also needs to be considered, potentially being single decisions (local), entire models (global) or the data. Local explanations could be helpful to base a decision on, whereas data attributions might be more helpful where the traceability of errors is relevant. While, strictly speaking, not part of the process of creating and showing explanations to users, it could also be a legal requirement to implement a way for end-users to object to decisions made by a model or to object to explanations which are conceived as false by the concerned user (Counterability of decisions/explanations). A further requirement can be Constancy, as it might be necessary to refer to a documented decision and its explanation to avoid liability. In addition to just documenting an XAI-explanation, in cases of ambiguous explanations (e.g., when multiple saliency maps are used to try to find general rules about how a computer-vison model decides) it might be necessary to not just document the explanation itself but accompany it by the interpretation of the person who used it to make a decision.
4. Legal Bases for the Use of XAI
When formulating legal requirements for the explainability of decisions made by ML-based applications, a stakeholder-based approach appears promising. This approach focuses on groups of people with a legitimate interest in XAI. Depending on the situation, different groups of people come into consideration, whereby four groups can be identified:
  1. (1)
    Persons who make a decision based on the recommendations/predictions of an ML model,
  2. (2)
    Persons whose data is processed by an ML application,
  3. (3)
    Persons who are responsible for the errors of an ML application and
  4. (4)
    Persons who want to assert claims against individuals responsible for the errors.
Group 2 stakeholders will have an interest in questioning the recommendations/prediction of the ML model, especially if it is disadvantageous to them. The interests of Group 3 stakeholders are similar to those of Group 1; here, too, liability should be avoided, but not for their own decision, but for errors in the ML application. Finally, Group 4 stakeholders seek to determine if the provider of the ML application can be held accountable for any errors. These definitions can be mapped to the definition of stakeholders for whom explanations are created described by Tomsett et al. (2018) as data-subjects and decision-subjects (group 1), operators and executors (group 2) and creators and examiners of ML systems (group 3 and 4). In the following, the legal situation of the respective stakeholders and the corresponding XAI-requirements are described.
4.1. ML Prediction as the Basis for a Human Decision: Plausibility Check
People sometimes make decisions based on ML recommendations/predictions. This is a challenge for various areas of law, such as criminal law (ML-supported sentencing), labor law (ML-supported personnel recruitment and management), administrative law, insurance law and others. The issues relating to company law are outlined below. However, it can be assumed that the legal problems in connection with ML-supported decision-making are comparable in all areas of law.
Company law deals with corporate fiduciaries (directors and officers of a company) who use ML applications when making decisions. From the legal point of view, the fiduciaries remain solely responsible for the decision made. If the decision is disadvantageous for the company, the fiduciaries are obliged to compensate the company for any resulting damages (fiduciary duty claims). A fiduciary can only avoid this liability if they prove that they did not breach any fiduciary duties (duties of care and loyalty) when basing a decision on the ML recommendation/prediction. Ways to achieve this are intensely discussed in legal literature. Many scholars consider even the use of black box ML applications by fiduciaries to be unproblematic (Petrin, 2019) while others argue that the lack of explainability leads the fiduciary to make an unfounded decision, thereby breaching the fiduciary duty of care (Cowger Jr., 2023). In Germany, the fiduciary duty of care has been specified by the Federal Supreme Court (BGH) in the so-called ISION judgement (BGH, 2011). The “ISION-Rule“ applies when the fiduciary seeks external advice due to a lack of in-house expertise. It has three requirements: (1) the expert consulted must be an independent professional who is qualified for the issue to be clarified, (2) the fiduciary must provide the expert with a comprehensive presentation of the company’s circumstances and disclose all necessary documents, (3) the fiduciary must subject the expert’s advice to a careful plausibility check. A breach of these requirements constitutes a breach of the corporate duty of care. Comparable regulations exist in various jurisdictions; for example, Section 141(e) of the Delaware General Corporation Law affords corporate directors a defense against liability only when the expert is chosen “with reasonable care”, and the director “reasonably believes” that the matter falls within the expert’s competence, relying in “good faith” on the expert’s advice.
In this context, the last requirement—the plausibility check—is particularly important. There is a consensus in the legal discourse that the above requirements should be applied to the use of an ML application (Buchholz and Dubovitskaya, 2023). However, it is disputed how the plausibility check should be carried out. Most authors seem to recognize that, at least in the case of black box models, sufficient traceability is not given and a plausibility check is therefore de facto impossible. For this reason, attempts are made to replace the plausibility check with less stringent requirements, such as requiring board members to familiarize themselves with the basic functions of AI applications and to regularly monitor the operation of the applications (Möslein, 2018). However, it remains unclear how such checks are to be carried out without the traceability of an AI decision. The requirements would therefore be reduced to the duty of fiduciaries to equip themselves with basic technical knowledge of ML applications. This does not enable effective and sufficient control of algorithmic decisions. Instead, decision-making authority shifts from human to machine. This remains true even if one resorts to the standard repertoire of organizational obligations, such as the duty to document decisions, appoint a Chief Information Officer (CIO), establish internal company rules for dealing with AI (e.g., through a special compliance brochure), and provide training for managers and other staff.
In order to be able to use ML models for fiduciary decisions, a plausibility check of their decisions needs to be possible and could be facilitated through XAI. To enable such a plausibility check, XAI systems need to provide certain properties, as discussed below. Note that properties, for which no relevant requirements are seen for this case, are left out.
Correctness is crucial in the use of XAI as explanations that do not align with the model’s decision do not offer a meaningful way to assess the plausibility of an ML decision. While explanations without Correctness may seem plausible, they can potentially increase automation bias (Vered et al., 2023) or hinder proper scrutiny of the decision system (Zhou and Joachims, 2023). To evaluate a decision accurately, all factors influencing the decision should be highlighted. To be able to reject a decision based on sensitive attributes like gender, a combination of Correctness and Completeness is necessary. Since most current XAI methods provide no guarantee for Correctness (except for limited axiomatic properties in SHAP), Consilience of the explanations of different methods could be used as a substitute. It is important to note that even when multiple XAI methods agree, Correctness of explanations cannot be guaranteed, just as human experts may make decisions based on reasons beyond their explanations. In the case of human expert advice, the fiduciary assesses the plausibility based on their understanding of the expert’s explanations. Applied to ML, this implies two criteria: explanations must be understandable to laypersons and to domain experts. This also requires Consistency, as an inconsistent explanation method can create confusion for non-experts who cannot resolve differing explanations and hyperparameter choices. Contrastivity is a necessary precondition to enable a user to discern between the prediction given by the model and relevant alternatives for action. For Covariate Complexity, the same as for Correctness can be stated: Ideally, the explanations should consider features and feature interactions as complex as in the original model, while still being understandable for laypersons. But given that simplifications are made explicit and the person tasked with the plausibility check still considers an explanation plausible, these requirements can be loosened. As described in Section 3.1, insufficient Compactness can render explanations unintelligible. On the other hand, explanations should be as complete as possible to highlight all relevant features. To resolve this conflict, each user should be shown the most complex explanation this user can still grasp, requiring an adaptation of the explanation to the Context. One way to give an XAI system this adaptability would be to make it controllable, allowing each user to adapt the explanation to their needs, e.g., by limiting the number of features highlighted in an explanation. Controllability can resolve such conflicts and additionally increase the amount of time a user spends with an explanation, increasing the explanation’s effectiveness (Cheng et al., 2019). Albeit Coherence to prior knowledge is not a requirement for explanations for plausibility checks, it will be a main part of the plausibility checks themselves, as laypersons and domain experts do not have other meaningful ways to assess ML decisions and explanations than comparing them to their prior knowledge. As the “ISION-Rule“ states, the expert consulted must be an independent professional qualified for the issue to be clarified. When extended to ML systems, the qualification of such systems for a general use case can be assessed via evaluation metrics such as accuracy, but the qualification for a specific data point with unknown label (as in real-world decisions) is more difficult to estimate. One main possibility to assess whether a model is qualified to make a decision for a specific scenario is the quantification of the Confidence of the model. Thus, a Confidence estimate of the model prediction should accompany the explanation for a specific decision. As explanations are not guaranteed to be correct, a Confidence estimate of the explanation itself is necessary to assess whether it can be relied upon. To facilitate the fiduciary’s defense in potential lawsuits, explanations should be presented in a way that allows for permanent storage and thus, Constancy. In the case of automatically generated ambiguous explanations, e.g., saliency maps, fiduciaries are advised to document their interpretation to prevent later disputes. Since the plausibility of a single decision needs to be assessed, local explanations should be sufficient. To improve the understanding of the limits of a ML system, it could be helpful to “calibrate one’s trust” via continued interaction with a model, requiring more than one local explanations (Ma et al., 2023). Legally, XAI could enable a comprehensive review of AI decisions, surpassing the level of the ISION principles. In legal discussions, there is debate about whether transparency requirements for algorithmic decisions should be no higher than those for human decisions. However, equal treatment of humans and machines is not required by law. On the contrary: if it is technically feasible to provide better explanations for algorithmic decisions compared to human decisions, there is no reason not to utilize this greater efficiency and consequently take it as a legal prerequisite.
Summarizing, when extending the ISION-principles to non-human explanations, they do not limit the type of explanation that can be used to assess the plausibility of a decision. While data attributions—showing similar decisions that happened in the past—could also be helpful here, it can be argued that such explanations are not specific enough to the decision at hand and require too much interpretation themselves. Thus, the simplest way to aid in a plausibility assessment is to use explanations such as feature importance with axiomatic properties for Correctness and Completeness such as SHAP, although the actual value of such properties is part of an ongoing discussion (Kumar et al., 2020; Merrick and Taly, 2019; Sundararajan and Najmi, 2020). As the requirement for Correctness is the most difficult one to fulfill, using multiple explanation methods (and thus, doing a “plausibility check” of the explanation itself) currently seems to be the best way to meet the legal requirements.
4.2. ML Prediction for the Assessment of Human Characteristics: Right to Explanation?
People are also increasingly becoming the recipients of algorithmic decisions, for example when ML is used to decide whether someone receives a loan or is invited to a job interview. As a result, decisions are being made on issues that are of central importance to social life. Especially if the decision is negative, the person concerned has an interest in finding out why. From a legal perspective, the question arises as to whether the person concerned has a right to an explanation of the decision and from which legal basis such a right can be derived. Of course, this raises the question of the extent of the legal guarantee, i.e., what must be disclosed/explained to the data subject.
It is being discussed whether the “right to an explanation” of algorithmic decisions can be derived from the provisions of the GDPR, specifically Article 15(1)(h) and Article 22(3) (Wachter et al., 2017; Wachter and Mittelstadt, 2018). In our legal opinion, this is indeed correct. Article 15(1)(h) grants the data subject the right to be informed of the existence of automated decision-making, including profiling, and to receive meaningful information about the logic, significance, and consequences of such processing. This right is further supported by Article 22(3), which allows the data subject to contest AI decisions. In the SCHUFA judgment, the Court of Justice of the European Union (CJEU) emphasized the need to interpret EU law not only of its wording but also of its context, objectives, and the purpose pursued by the act of which it forms a part (CJEU, 2023). Given the GDPR’s goal of enabling effective exercise of data subjects’ rights, the “right to an explanation” is already inherent in the GDPR. The Advocate General at the CJEU Priit Pikamäe, in his opinion on the SCHUFA case, shares a similar sentiment. He considers that “the obligation to provide ’meaningful information about the logic involved’ must be understood to include sufficiently detailed explanations of the method used to calculate the score and the reasons for a certain result. In general, the controller should provide the data subject with general information, notably on factors taken into account for the decision-making process and on their respective weight on an aggregate level, which is also useful for him or her to challenge any ‘decision’ within the meaning of Article 22(1) of the GDPR.”(Advocate General Pikamäe, 2023) Besides the GDPR, the right to explanation of algorithmic decisions may find grounding in fundamental rights. At the EU level, considerations of human dignity and the right to the protection of personal data (Articles 1 and 8 of the EU Charter of Fundamental Rights) come into play. Article 8(2) sentence 2 of the EU Charter of Fundamental Rights affirms every person’s right to access data collected about them and the right to have it rectified. However, it is questionable whether this guarantee also includes the right to explanation of decisions that are the result of ML-based data processing. The narrow formulation of Article 8(2) sentence 2 may not explicitly encompass such rights. Nevertheless, ML-based data processing, like any data processing, constitutes an interference with the right to the protection of personal data, necessitating justification. The proportionality test allows for weighing the data subject’s interest in decision explanation against the ML application operator’s interest in protecting trade secrets (Article 17(2) of the EU Charter of Fundamental Rights, protection of intellectual property). The principle of practical concordance mandates finding a balance that is as gentle as possible, ensuring neither conflicting fundamental right takes precedence over the other (akin to Pareto optimum in game theory). Thus, the goal is to explain the ML prediction to the concerned individual without revealing algorithmic components constituting trade secrets. However, the challenge of reverse engineering, often permitted in some national laws (e.g., in the German Law on the Protection of Trade Secrets), must be addressed, e.g. by new regulations.
The task for XAI regarding a “right to explanation” (independent of the legal basis used) can be summarized as: Enable end-users of AI systems to object to ML decisions and—as argued in (Wachter et al., 2018)—show ways to achieve a different decision. As these tasks require local explanations as in Section 4.1 with mainly the same requirements—a plausibility check of the decision by a layperson—only the differences are discussed here.
The GDPR grants users the right to object to a decision in Article 22(3) (here called Counterability), thus from a practical standpoint, a process for objecting automated decisions should be implemented. This could also entail the option to object to explanations which are deemed unjust or incorrect, especially since Correctness cannot be guaranteed. As much as a Confidence estimate of a decision and its explanation could help users to assess whether a model was certain in their case or whether it was more likely to be wrong, it could also complicate the privacy of such systems by enabling membership inference attacks (Shokri et al., 2016). To prevent issues such as this, other measures could be taken, e.g., limiting the amount of explanation requests a user can make. Similar problems arise for data influence explanations, as sensitive training data (other than their own) should not be shown to end-users. Compared to fiduciary decisions, explanations can be generated when need arises, as it is not feasible (nor necessary) to compute explanations for all possible users in advance. In our opinion, merely providing counterfactual explanations, as recommended in (Wachter et al., 2018), may not be sufficient to achieve the goal of enabling users to object to ML decisions and give them the opportunity to alter the outcome. This viewpoint is underscored by the statement of Advocate General Priit Pikamäe at the CJEU (Advocate General Pikamäe, 2023). Because counterfactual explanations do only provide a very limited account of how a decision was made (“This feature must change to change the outcome, thus it is most likely important.”), further information about the decision needs to be given, which could for example take the form of feature importance explanations as in Section 4.1. Thus, to fulfill all aims of the use of XAI regarding the ‘right to explanation’, several types of explanations could therefore be jointly given to users.
4.3. ML Defects: Product Safety, Product Liability and Contractual Liability for Defects
XAI-methods are often stated to help in debugging models or showing the fairness of such systems (Deck et al., 2023; Langer et al., 2021), both relevant aims in product safety, product liability and contractual liability for defects. In the following, these areas and their relation to XAI are described, followed by a discussion of properties XAI-methods need to fulfill to truly aid in these cases.
4.3.1. Product Safety
Under certain conditions, developers of ML applications must comply with the requirements of the Product Safety Law. The European Union’s regulations on product safety are outlined in the General Product Safety Directive (GPSD) of December 3, 2001, which is set to be superseded by the new General Product Safety Regulation (GPSR). The GPSR came into effect on June 12, 2023, and its application will commence on December 13, 2024. In Germany, product safety for non-food products is governed by the German Product Safety Act (ProdSG). It incorporates the regulations of European Product Safety Law and applies when products are introduced to the market, distributed, or used for the first time as part of a business activity (Section 1(1) of the Product Safety Act). A product is a good, a substance or a mixture that has been manufactured through a production process (Section 2 no. 21 of the Product Safety Act). According to the prevailing view, a product requires a physical object, so that stand-alone software is not a product within the meaning of this act. Thus, it only becomes applicable to software when installed on data carriers or embedded in a physical product. In this case, at least the general requirement applies that the product must not endanger the safety and health of persons when used as intended or in a foreseeable manner (Section 3(2) sentence 1 of the Product Safety Act); if the product is subject to one or more special legal regulations mentioned in Section 8 of the Product Safety Act, it must also meet the requirements specified in these legal regulations. The person who places the product on the national market is therefore obliged to ensure that the product is sufficiently safe to exclude risks to the physical safety and health of persons, with measures adapted to specific hazards. However, absolute safety is not owed; rather, the manufacturer must (only) take the measures that are objectively necessary and reasonable according to objective standards in the circumstances of the specific case.
If an ML application is used in the manufacture of a product without becoming part of this product (e.g., for quality control in product manufacturing processes), the obligations under the Product Safety Act only apply to the product manufacturer (supplier). A fault in the ML application can potentially lead to a safety-relevant fault in the end product. In such a case, the manufacturer will have an interest in recourse against the person who has contractually provided them with the ML application. Even without a special agreement, such a possibility of recourse will regularly arise from the contractual obligation between the manufacturer and the supplier (Section 280(1) of the German Civil Code). Product manufacturers will likely seek to negotiate contractual guarantees and liability clauses with their suppliers. Then the supplier, who is the developer of the ML application, would have an increased interest in ensuring that their ML application is free of defects (given they are the manufacturer) or in agreeing contractual guarantees with the supplier of the application.
4.3.2. Product Liability
Product liability in the European Union is currently governed by the (almost 40-year old) Directive 85/374/EEC on Liability for Defective Products, implemented in Germany through the Product Liability Act (ProdHaftG). Civil liability (liability for damages) under Section 1(1) of the Product Liability Act applies if a product defect causes death, personal injury, or damage to property other than the defective product. This requires that the product is intended for private use or consumption and has been used accordingly. The current interpretation of “products” includes movable objects and electricity (Article 2 Product Liability Directive; Section 2 ProdHaftG), but it is unclear if non-physical software is covered. German literature tends to include software under product liability (as in the proposed new Product Liability Directive).
Section 2(2) no. 2 of the Product Liability Act states that the manufacturer must ensure that the product is free of defects when placed on the market. According to Section 2(2) no. 5 of the Product Liability Act, the manufacturer’s liability is excluded if the defect could not have been recognized according to the state of the art in science and technology at the time when the manufacturer placed the product on the market (so-called development risk). However, the fact that the manufacturer of an ML application cannot foresee its behavior in a specific hazardous situation does not constitute a development risk. The blanket exoneration of manufacturers of ML systems is therefore out of the question.
The proposed directive extends the requirements for freedom from defects: According to Article 6(1)f, a product is also defective if it does not fulfill the security-relevant cybersecurity requirements. At the same time, the proposed directive specifies the requirements for the manufacturer. The manufacturer cannot be held liable if they were unable to detect the defect based on objective scientific and technical knowledge at the time of sale or during their control of the product. However, the manufacturer must ensure that both the software and products containing software are defect-free throughout the duration of their control. During this period, the manufacturer must also provide the updates and upgrades necessary to maintain safety (Article 10(2) of the proposed directive). “Control of the manufacturer” means the fact that the manufacturer of a product authorizes: a) the integration, connection or supply of a component, including software updates or upgrades, by a third party or b) the modification of the product (Article 4(5) of the proposed directive).
In addition to the Product Liability Act, there is also tortious product liability in Germany in accordance with Section 823(1) of the German Civil Code. Here, the manufacturer is liable for damages to anyone whose legal interests (in particular life, body, health and property) have been injured by a defective product. This presupposes that the manufacturer has breached marketing obligations (design, manufacturing, instruction, product monitoring, warning and recall obligations). In the context of Section 823(1) of the German Civil Code, it is recognized that the manufacturer is obliged to constantly adapt the software to new knowledge about safety risks and to exploit new technical possibilities. In contrast, the Product Liability Act does not recognize obligations to monitor products, issue subsequent warnings or recall defective products.
4.3.3. Liability for Defects (Seller’s Point of View)
The manufacturer or supplier of an ML application is liable to its contractual partner (client, buyer, etc.) for defects in accordance with the provisions of the German Civil Code on contractual liability for defects (warranty law). A defect exists if the subject matter of the contract does not fulfill the contractual requirements, considering the objective or average requirements in the absence of specific contractual agreements. Therefore, liability extends beyond damages to health or property, as in product liability law, and covers any deviation from the contractually agreed quality. On the other hand, liability for defects is narrower than product liability, as it applies only to the contractual partner. For contractual liability, it does not matter whether the ML application is integrated into a physical object or whether it is stand-alone software.
Which provisions are relevant in the individual case depends on the respective type of contract (purchase, contract for work, lease, etc.). In a purchase agreement, for example, the seller’s liability for defects under Section 437 of the German Civil Code is applicable. In this case, the buyer is entitled to special warranty rights such as the right to subsequent fulfillment, withdrawal or reduction regardless of whether the supplier of the ML application is responsible for its defectiveness. The buyer’s claims for damages, however, require fault on the part of the seller. In the absence of such special warranty provisions, liability arises from the general standard of Section 280(1) of the German Civil Code, because the defect in the ML application constitutes a breach of the obligations arising from the contract. However, this general provision governs liability for damages and therefore presupposes culpability (fault). The parties are free to agree on stricter or more lenient liability terms. Special provisions apply if the supplier of the ML application is an entrepreneur and its contractual partner is a consumer. In this case, deviation from statutory provisions (Sections 327 et seq., 475b et seq. of the German Civil Code) is only permitted under strict conditions (Sections 327h, 327s, 476 of the German Civil Code).
The above three Sections 4.3.1 to 4.3.3 result in the same task for XAI: Find and if possible rule out defects in the ML model. Different solutions can be devised, ranging from global explainability of a model to rule out defects, to using local explanations or data attributions to find defects, to possibly not using XAI methods at all given sufficient testing and documentation. Ways in which XAI can assist to find or rule out defects are covered in the following, for a way to ensure safety by documentation see (Brajovic et al., 2023). In all of these cases, explanations can be expected to be generated for expert users.
To ensure freedom of defects, Correctness is at the heart of XAI use, as a global explanation method that does not conform to the model might miss out on defects. To date, no such global explanation method seems to be known and computationally feasible, as highlighted in the closely related field of formal verification of neural networks (Albarghouthi, 2021). The only way to provide an explanation with such properties is to use explainable/interpretable models, as demanded by Rudin (2018) for safety critical use cases. To spot defects within black-box AI applications, local explanations can help, where Correctness-problems could once again be mitigated by using the Consilience of multiple XAI-methods, although this would still not give providers of such system enough information to guarantee freedom of defects. Since XAI in the near future seems more likely to be able to show the presence of defects than to provide correct global explanation methods on which a guarantee for the freedom of defects could be built, the first case will be considered for the following properties.
As long as the explanations do show relevant defects, there is no need for Complete explanations, as for instance even a very specific method which only highlights whether a protected attribute is used in a decision could help to spot unwanted behavior. The same can be argued for Compactness, as not entire explanations need to be understood but they can be searched for parts where defects are suspected. In practice, it could prove difficult to choose the (number of) features for such explanations. In the search for defects, inconsistent explanation methods can be used, as running them multiple times seems feasible to potentially spot defects. Additionally, they do not necessarily need to provide Continuity, unless a global understanding of a model should be gained, which is only possible if explanations are continuous and can thus be generalized to near data points. Since defects, for which liability could become relevant, e.g., discriminating decisions, can also result from the interaction of input features for instance via confounding variables, used XAI-methods should provide the maximum understandable amount of Covariate Complexity. Additional to using XAI, confounders could be discovered via methods as described by Rogozhnikov et al. (2022) to ensure that no sensitive attributes are reconstructed based on other input features (Wachter and Mittelstadt, 2018). A Confidence estimate could itself show unwanted behavior in a model, e.g., a model being uncertain about a safety-critical decision which is expected to arise during its deployment. Because of this, model Confidence should be considered. Since the Controllability of an explanation improves the user’s understanding of the underlying model and thus, the chance of finding relevant defects, explanations should be interactive. To avoid liability claims and prove a freedom of defects (or the safety of the ML system), the relevant explanation artifacts and the way in which they were created need to provide Constancy.
Overall, there is no clear consensus on how XAI can help to spot or rule out defects in ML applications. While a human intelligible and fully correct global explanation seems ideal, it also seems unrealistic for some use cases as of now—partly due to the currently available ML and XAI methods, but also due to the complexity of some use cases. The closest to this ideal are interpretable models, for which guarantees can be derived, given that they reason on human-intelligible features. If no such model can be devised for a certain use case, the best way forward seems to be a joint combination of documentation and a variety of explanation methods for the model itself, its explanations and the data used.
4.3.4. Liability for Defects (Buyer’s Point of View)
The buyer who receives a defective product and intends to take action against the seller under Section 437 of the German Civil Code must prove that the product was defective at the time of the transfer of risk, generally at the time of delivery. Such a defect may be presumed only if the buyer is a consumer (Section 477 of the German Civil Code) Furthermore, the buyer may have an interest in determining whether the seller is responsible for the defect, for instance, if the defect could have been discovered during production. This information could be necessary when the buyer asserts claims for damages.
This tasks XAI with showing the existence of a defect and finding the cause of it, as this way the liability of the provider of an AI system (or even the data provider) can be proven. This can be expected to be done by AI experts. To find (reasons for) defects, all parts of an AI application can be inspected—as in the use case above with local and global explanations. A special focus for the liability from the buyer’s point of view lies on data attribution, as for example local explanations can show the presence of erroneous behavior, but not its cause, which is expected to originate in the training data used or in the development of the model itself. While the second case should be checked via documentation as proposed in (Brajovic et al., 2023), the first one—attributing certain behavior to training data samples—can be solved via data attribution.
To make sure that only true reasons for a defect are shown and thus, no false liability claims are made, Correctness of explanations in combination with Completeness is necessary to not miss out on causes of defects. Contrastivity is necessary to be able to discern between different causes for defects, otherwise the users of such systems might be burdened with the interpretation of which possible cause for a defect was relevant for specific data points, rendering the use of XAI pointless. As in the previous use cases, Covariate Complexity needs to be considered in the explanations, as the interactions of features can themselves highlight relevant causes for defects. Since Confidence estimates can point towards problems such as unbalanced datasets—which can lead to defects themselves—the model Confidence should be considered for the ML model. Like in the previous cases, Controllability of explanations helps to better understand the ML model and its explanations, and thus, should be utilized.
Overall, the current state of the art of XAI is only partly suitable for identifying the causes of defects, as most current explanations provide rather superficial answers and require user interpretation to find the cause for certain behavior. This can easily be described for saliency explanations: While they can show whether or not something has been detected in an image, they do not provide any information about why this has happened. The most promising direction towards finding defect causes via XAI seem to be data attribution explanations, where defects can be traced down to (possibly erroneous) data samples. A further way to investigate such defects could be causal attribution mechanisms in addition to XAI.
5. Discussion
While the EU AI Act is seen as a breakthrough in AI regulation, it is important to note that it is not expected to supersede all other relevant laws influencing the use of AI (and thus, XAI). The laws described in Chapter 4, from which the requirements were derived, mainly pertain to civil liability and will coexist with the AI Act, which falls under public law. Although these laws interact (unfulfilled requirements of the AI Act might be seen as defects in the sense Sections 4.3.3 and 4.3.4), any further discussion on their possible interactions is beyond the scope of this paper.
There has been much debate about the perceived conflict between regulation and innovation in the context of the AI Act. However, it is important to note that regulation itself does not stifle innovation; rather, unclear regulation or instructions may impede progress. To avoid such ambiguity in the field of XAI, this paper proposes a structured approach based on specific properties to guide XAI use and development. Similarly, international or national norms could recommend XAI properties or specific methods for specific use cases, providing clear instructions on which methods to employ under different circumstances. These norms should consider additional legal foundations for XAI, such as criminal law, administrative law, and insurance law.
To promote advancements in the intersection of law and XAI, technical and legal experts should actively exchange information. It is crucial to bridge the gap between the two fields, as their current understanding of XAI differs significantly. While legal experts often emphasize the need for transparency to ensure safe AI usage and address liability issues, the current technical possibilities do not fully support such a broad statement. Additionally, the terminology used in both fields may not align, leading to communication challenges (Schneeberger et al., 2023). While the collaboration and integration of both fields are important, it is worth noting that the law cannot fully dictate the use of technology. In the case of XAI, other disciplines such as social sciences (examining the effectiveness of explanations) and user interface design (determining how explanations should be presented for desired effects) can contribute to filling gaps in legal requirements. As the law provides minimum requirements, insights from these fields become more crucial for properties not touched by legal requirements (such as Compositionality and others) and when surpassing these minimum standards.
6. Summary
In this paper, the relationship between law and XAI was explored, deriving requirements for different areas of law based on an extended taxonomy of XAI properties. The areas of law considered were fiduciary decisions, GDPR, and product safety and liability. These areas require different XAI properties and methods. For fiduciary decisions, assessing plausibility is important, similar to the GDPR, where users should be able to object to decisions and receive actionable advice. In product safety, identifying defects in ML models is crucial, while product liability focuses on demonstrating the causes of defects. A tabular summary of the derived requirements and explanation recommendations can be found in Table 1 as a guidance for working with XAI in these areas.
Various potential aims for XAI have been proposed in the literature (see for example Saeed and Omlin (2023)), but the practical utilization of XAI to achieve these goals remains unclear. The lack of consensus regarding evaluation metrics for the objectives and properties of XAI further complicates the matter, as the reliable and valid assessment of XAI method characteristics is a necessary precondition to meaningfully confirm or reject specific methods for a given use case. Possible actions to address this could involve developing new evaluation metrics or conducting studies to assess the effectiveness of XAI in specific settings. However, while the legal practice often requires the use of state of the art technologies, it appears that the current state of XAI is not fully equipped to fulfill all legal requirements. The most challenging XAI properties in this regard are Correctness (explanations conforming to model logic), Confidence estimates of models and explanations themselves, and the lack of global explanation methods, indicating the need for further research in these areas. One ongoing study by some of the authors focuses on the use of XAI in science and industry. It remains to be seen if practitioners in these areas face similar challenges. To inform future development, law can be used as an inspiration for XAI-methods with specific properties or new ways to evaluate XAI. To this end, we urge practitioners from both fields to collaborate, so that a common terminology and a common understanding of the capabilities of XAI can be established.
The requirements for XAI presented in this paper serve as recommendations that contribute to the ongoing discussion on the utilization und future development of XAI. We eagerly anticipate the future advancements in this field, both from a technical standpoint and within the legal and normative realms.
Table 1. Requirements for XAI derived from the different legal areas. Checkmarks denote that a property is necessary for that specific use case. (1) denotes that instead of Correctness, Consilience of explanations can be used. (2) denotes the Tradeoff between Completeness and Compactness, (3) between Completeness/Covariate Complexity and understandability, (4) between Confidence estimates and privacy. (5) denotes that multiple predictions could be used to ”calibrate one’s trust”, e.g. to learn when and when not to trust the ML model.
Legal Basis Plausibility Check - Section 4.1 Right to Explanation - Section 4.2 Safety/Liability - Sections 4.3.1-4.3.3 Liability (Buyer) - Section 4.3.4
XAI-Task Check the ML-decision for plausibility Enable user to understand and object to decision Show the absence of errors (as far as possible) Show the cause of an error
Audience of Explanation Laypeople, domain experts Laypeople ML experts ML experts
[…]
Consistency $\checkmark$ $\checkmark$
Continuity
Contrastivity $\checkmark$ $\checkmark$ $\checkmark$
[…]
Compositionality
Confidence $\checkmark$ (for prediction and explanation) $\checkmark$ (for prediction and explanation) (4) $\checkmark$ (for prediction) $\checkmark$ (for prediction)
Context $\checkmark$ $\checkmark$
Coherence
Controllability $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
[…]
Computations
Coverage Prediction (5), Data Prediction Model, Prediction, Data Model, Prediction, Data
Counterability $\checkmark$
Constancy $\checkmark$ $\checkmark$ $\checkmark$
Recommendation for Explanation Type Feature Importance (or similar) Counterfactuals and Feature Importance (or similar) Global/interpretable models Data attribution
Acknowledgements.
Parts of this paper were created with the help of a data-centric implementation of ChatGPT (3.5 turbo). It was used to create LaTeX-code for tables and figures and to refine the drafts of some sections. This paper is funded in parts by the German Federal Ministry for Economic Affairs and Climate Action under grant no. 19A21040B (project ”veoPipe”).
Appendix A Summary of used XAI Properties
A short description of all used XAI properties is provided in Table 2. A more detailed description can be found in Chapter 3.
Table 2. 12 Co-Properties of Nauta et al. (2023) and additional Process-Properties (table adopted from (Nauta et al., 2023)).
Property Definition
Correctness Describes how faithful the explanation is w.r.t. the black box.
Completeness Describes how much of the black box behavior is described in the explanation.
Consistency Describes how deterministic and implementation-invariant the explanation method is.
Continuity Describes how continuous and generalizable the explanation function is.
Contrastivity Describes how discriminative the explanation is w.r.t. other events or targets.
Covariate Complexity Describes how complex the (interactions of) features in the explanation are.
Compactness Describes the size of the explanation.
Compositionality Describes the format and organization of the explanation.
Confidence Describes the presence and accuracy of probability information in the explanation.
Context Describes how relevant the explanation is to the user and their needs.
Coherence Describes how accordant the explanation is with prior knowledge and beliefs.
Controllability Describes how interactive or controllable an explanation is for a user.
Consilience Describes whether more than one XAI-method should be used to fulfill the legal requirements.
Computations Describes when an explanation needs to be available, e.g. right when a prediction is shown or at a later point of time.
Coverage Describes what should be explained, a single prediction (local), an entire model (global) or the influence of training data samples (for predictions or the overall model behavior).
Counterability Describes whether a process needs to be implemented which allows end-users to object to AI decisions or explanations.
Constancy Describes whether explanations need to be presented in a format that can be saved for later examination.